OpenStudy (anonymous):

Simplify -15w^0u^-1 ----------- 5u^3

OpenStudy (anonymous):

click good answer

OpenStudy (anonymous):

\[-15w ^{0}u ^{-1}\div5u ^{3}\]

OpenStudy (diyadiya):

ok so first of all \(w^0=1\)

OpenStudy (diyadiya):

so now we have \[\frac{-15u^{-1}}{5u^3}\]

OpenStudy (diyadiya):

Now we have this formula \[\large \frac{x^m }{x^n}=x^{m-n}\]

OpenStudy (diyadiya):

Can you solve now ?

OpenStudy (diyadiya):

@hogie

OpenStudy (diyadiya):

\[\large \frac{-15u^{-1}}{5u^3}= \frac{-3u^{-1}}{u^3}= -3u^{(-1-3)}\]